OpenStudy (anonymous):

find the x int for 4x^2+24x-24

OpenStudy (anonymous):

4x^2+24x-24 = 0 [set equation equal to zero] x^2+6x-6 = 0 [divide both sides by 4] a = 1, b = 6, c = -6 Use quadratic formula

OpenStudy (anonymous):

\[x = -3 \pm \sqrt{15}\]

OpenStudy (anonymous):

so if you cannot factor the middle term you have to use quadratic?

OpenStudy (anonymous):

Yes, if you cannot factor, you must use the quadratic formula (or complete the square).

OpenStudy (anonymous):

thanks